# Trajectory Volatility for Out-of-Distribution Detection in Mathematical Reasoning

### v1

@@ 4.4 Beyond Mathematical Reasoning
Apart from mathematical reasoning, our method also has a wider range of potential applications that can be extended to any task where the output space exhibits the pattern collapse property. An example would be multiple-choice questions, which is a popular evaluation tool in the era of large language models and also display the pattern collapse property due to the limited output space being confined to the “ABCD” four options. To verify the generalizability of our method, we conduct experiments using the multiple-choice dataset MMLU [12], and our method also outperforms all traditional algorithms in this setting. Results and analyses are shown in Appendix F.
5 Discussion
Hyperparameter: Smoothing Order $k$.
In the main experiments, we find that the performance of differential smoothing fluctuates greatly in different settings. Therefore, we conduct the ablation of the smoothing order $k$. Results and analyses are shown in Appendix G.1.
Dilemmas of Embedding Representations in Reasoning.
Figure 4: A toy embedding projection. $\mathrm{A}$ is the benchmark, $\mathrm{X}$ is the ID sample, $\mathrm{Y}$ and $\mathrm{Z}$ are OOD samples.
We explore why the input-embedding-based method is not powerful under mathematical reasoning, as prior work [46] has shown it to be very effective in text generation scenarios.
[…]
We also use powerful sentence embedding techniques like SimCSE [11] to verify this conclusion. Results and analyses are shown in Appendix G.2.
Can Chain-of-Thought Address Pattern Collapse?
An intuitive approach to addressing pattern collapse in the output space is to use chain-of-thought (CoT) techniques [54, 20] to expand the output space size. Therefore, we include the solution steps in the output to re-evaluate the output embedding results and find that CoT cannot address this issue. Results and analyses are shown in Appendix G.3.
6 Related Work
@@
7 Conclusion
Limitations
In this part, we discuss the limitations of this paper and outline our efforts to make up for them, along with potential avenues for future exploration.
  • Dataset Size: Due to the difficulty of collecting and labeling mathematical reasoning data, dataset sizes in this field are generally small, mostly in the hundreds or thousands, making it difficult to obtain millions of data for training and reasoning as in translation or summarization tasks. To address this, we adopt test sampling to reduce the randomness under small-scale testing and mitigate the data imbalance.
  • Generalizability: Our motivation and excavated phenomena are the mathematical reasoning scenario, and thus our main experiments are mainly centered on mathematical reasoning. However, the motivation for the method is primarily related to pattern collapse in the output space, and thus our method can be theoretically extended to any task where the output space has this property, such as multiple-choice questions. We have provided a preliminary demonstration of the extension feasibility in Appendix F, and more exploration can be left to future work.
Ethics Statement
Appendix A Motivation I: Theoretical Analysis
@@ A.1 Problem Setup
For a given sample $s_{i}$, $[{\boldsymbol{y}_{i}}]_{l}\in\mathbb{R}^{d}(1\leq l\leq L)$ denote the hidden embedding of the $l$-th layer. We define the embedding coordinate in the latent space as $\{(l,[{\boldsymbol{y}_{i}}]_{l})\}_{l=1}^{L}$.
@@
We assume the ${\boldsymbol{F}}_{i}(x)=\left[F_{i1}(x),F_{i2}(x),...,F_{id}(x)\right]^{\top}:\mathbb{R}\rightarrow\mathbb{R}^{d}$ with $d$ independent components fits the $L$ coordinates, representing a continuous learning trajectory for $s_{i}$, so $[{\boldsymbol{y}_{i}}]_{l}={\boldsymbol{F}}_{i}(l)$. ${\boldsymbol{F}}_{i}(x)$ is taken from the functional space $\boldsymbol{X}$, so $\{{\boldsymbol{F}}_{i}(l)\}_{l=1}^{L}$ are all independent variables. We constrain each component function of this ${\boldsymbol{F}}_{i}(x)$ to satisfy the $m$-order $(m\geq 3)$ derivability property. Under this setting, the definition of component-independent trajectory volatility (Eq. 3) equates to
${\boldsymbol{V}}(s_{i})$ $=\left[V_{1}(s_{i}),V_{2}(s_{i}),...,V_{L}(s_{i})\right]^{\top}$ (13)
@@
We specify that $[{\boldsymbol{y}}_{i}]_{L}={\boldsymbol{F}}_{i}(L)\sim\mathcal{N}(\boldsymbol{c},\Sigma^{2})$, where
$\boldsymbol{c}=[c_{1},c_{2},...,c_{d}]^{\top},~{}~{}~{}\Sigma=\mathrm{diag}(\delta_{1},\delta_{2},\cdots,\delta_{d}).$ (14)
According to the pattern collapse property under different tasks, we can constraint the endpoint embedding ${\boldsymbol{F}}_{i}(L)$ in the output space:
  • For mathematical reasoning with pattern collapse, $\Sigma\rightarrow\boldsymbol{O}$, so we approximate that ${\boldsymbol{F}}_{i}(L)\equiv\boldsymbol{c}$;
[…]
@@
Theorem A.1 (Main Theorem)
@@
$\mathbb{E}_{\{{\boldsymbol{F}}_{i}(l)\}_{l=1}^{L},~{}\{{\boldsymbol{F}}_{j}(l)\}_{l=1}^{L}~{}\sim~{}\mathcal{U}(\mathbb{R}^{d})}\left\{{\boldsymbol{V}}(s_{i})-{\boldsymbol{V}}(s_{j})\neq{\boldsymbol{0}}|{\boldsymbol{F}}_{i}(L),{\boldsymbol{F}}_{j}(L)\equiv\boldsymbol{c}\right\}$
@@ A.2 Prelinimary
Proposition A.1 (Lagrange Remainder Term)
@@
Lemma A.1 (Error Bound for the Midpoint Rule)
@@
Following Theorem A.1, we do the inequality scaling as below:
@@
Lemma A.2 (Differential-Integral Error Order Estimation)
@@
Following Lemma A.1, we do the inequality scaling for the first-order derivative summation:
@@
A.3 Main Theorem Proof
Finally, we formally prove our main theorem A.1 in this sub-section, thereby verifying the Hypotheses 1 in the main paper. Since the $d$ components of ${\boldsymbol{V}}(s_{i})-{\boldsymbol{V}}(s_{j})$ are independent of each other, we only consider the $d$-th dimensional component and the rest of the components $V_{d}(s_{i})-V_{d}(s_{j})$ can be proved in the same way.
Proof.  According to Lemma A.2, we can approximate $V_{d}(s_{i})-V_{d}(s_{j})$ as
@@ A.3 Main Theorem Proof
We assume that the first interval $(1,i_{1})$ and $(1,j_{1})$ are constant positive intervals (the same can be proven for constant negative intervals). In this setting, we can continue to simplify Eq.(26):
@@
$\boldsymbol{c}=\left[(-1)^{i_{p}-1}\cdot F_{id}(L)-(-1)^{j_{q}-1}\cdot F_{jd}(L)\right],$ (30)
then we can rewrite Eq.(29) to the matrix form:
$\int_{1}^{L}\left|\frac{\mathrm{d}F_{id}}{\mathrm{d}x}\right|~{}\mathrm{d}x-\int_{1}^{L}\left|\frac{\mathrm{d}F_{jd}}{\mathrm{d}x}\right|~{}\mathrm{d}x=\boldsymbol{A}\boldsymbol{x}+\boldsymbol{c},$ (31)
where $\boldsymbol{A}$ is the coefficient matrix, and $\boldsymbol{x}$ is the unknown variable vector:
$\boldsymbol{A}=[-1,1,\underbrace{2,-2,2,-2,...,(-1)^{p-1}\cdot 2}_{p~{}\text{dimensions}},\underbrace{-2,2,-2,2,...,(-1)^{q}\cdot 2}_{q~{}\text{dimensions}}]\in\mathbb{R}^{1\times(p+q+2)}$ (32)
$\boldsymbol{x}=[F_{id}(1),F_{jd}(1),F_{id}(i_{1}),...,F_{id}(i_{p}),F_{jd}(j_{1}),...,F_{jd}(j_{q})]^{\top}\in\mathbb{R}^{(p+q+2)\times 1}$
We let $i_{p}\equiv j_{q}(\mathrm{mod}~{}2)$, and return to the condition in the mathematical expectation on both sides of the inequality in the main theorem A.1 for the following categorical discussion:
  • If $F_{id}(L)=F_{jd}(L)=c_{d}$, then $\boldsymbol{c}=\boldsymbol{0}$, so $V_{d}(s_{i})-V_{d}(s_{j})=0\Longrightarrow\boldsymbol{Ax}=\boldsymbol{0}$;
[…]
We denote $N(A)$ as the solution space of $\{\boldsymbol{x}|\boldsymbol{Ax}=\boldsymbol{0}\}$ and $P(A)$ as the solution space of $\{\boldsymbol{x}|\boldsymbol{Ax}=-\boldsymbol{c}\}$. For $x\in P(A)$, $x=x_{p}$ or $x=x_{p}+x_{n}$ where $x_{p}$ is the special solution and $x_{n}$ is the solution in zero space, i.e., $x_{n}\in N(A)$. Since $\mathrm{rank}(A)=1<p+q+2$, the special solution $x_{p}$ exists, so the size of the solution space $P(A)$ is larger than $N(A)$. In this case, when the variable $\boldsymbol{x}$ is sampled from the real number space, its probability of being in $N(A)$ is smaller, namely its probability of not being in $N(A)$ is larger. This is exactly equivalent to the form of the proof of Main Theorem A.1, so the proof is complete. $\square$
A.4 Extended Conclusion

### v2

@@ 5.2 Beyond Mathematical Reasoning
Apart from mathematical reasoning, our method also has a wider range of potential applications that can be extended to any task where the output space exhibits the pattern collapse property. An example would be multiple-choice questions, which is a popular evaluation tool in the era of large language models and also display the pattern collapse property due to the limited output space being confined to the “ABCD” four options. To verify the generalizability of our method, we conduct experiments using the multiple-choice dataset MMLU [10], and our method also outperforms all traditional algorithms in this setting. Results and analyses are shown in Appendix F.4.
6 Rethink Inapplicability of Static Embedding Methods
In Figure 1, we find that the static embedding space for both input and output fails to capture distinct features across different domains. This renders traditional embedding-based methods unsuitable for mathematical reasoning tasks. Experimental results in Section 4 also verify this disadvantage. In this section, we conduct a detailed analysis of these phenomena, and reveal the root cause of the poor performance of static embedding methods in mathematical reasoning scenarios.
6.1 Input Space: Representation Dilemma on Mathematical Expressions
The input space in mathematical reasoning exhibits vague clustering features compared to text generation scenarios as shown in Figure 1. We speculate that as semantic representations, embeddings cannot accurately measure mathematical expressions in the mathematical sense.
Table 4: A toy example about cosine similarities between different mathematical expressions.
Mathematical Expression Cosine with Benchmark
Benchmark: 2+3+4=?
8-0+9=? 0.65
x^2*y^3*x^4=? 0.83
234*2+345*4+456*4-243*3=? 0.78
We begin by considering a counter-intuitive toy example. We construct a benchmark expression “A: 2+3+4=?” and three diverse expressions: “X: 8-0+9=?”, “Y: x^2*y^3*x^4=?”, and “Z: 234*2+345*4+456*4-243*3=?”. From the mathematical sense, X is ID for A, whereas Y and Z are OOD for A since the difficulty and knowledge used to solve Y and Z are largely different from A compared to X. Table 4 presents the cosine similarity between each of them and A. Notably, X exhibits the lowest similarity to A, primarily due to having fewer token sharing with A. This is inconsistent with human perception of mathematics.
OOD Domain
(a) Input Embedding MaDis (b) Output Embedding (w/ CoT) MaDis
ID Domain algebra - 63.77 80.01 58.79 50.80 - 71.37 84.24 65.06 56.71
geometry 85.68 - 88.14 86.55 69.03 72.12 - 92.42 86.08 70.89
cnt.&prob 49.95 44.94 - 38.02 51.76 46.19 53.61 - 40.15 51.63
num.theory 66.35 78.87 50.00 - 71.93 63.02 78.76 69.85 - 67.96
precalculus 85.14 79.31 86.18 89.34 - 79.84 85.65 86.77 90.35 -
To demonstrate this phenomenon, we use one of five domains from the MATH dataset as the ID and the remaining four domains as OOD. We rotate this setting across five iterations. Under each ID-OOD pair, we allocate half of the ID samples to compute the ID embedding distribution, while the other half and all OOD samples constitute the test samples. The Mahalanobis distance is then calculated between embeddings and the ID distributions for each sample input, yielding the OOD score. We employ SimCSE [9] to generate sentence embeddings. Table 5(a) presents the AUROC score matrix, in nearly half of the settings, the AUROC value hovers around 50, i.e., completely random. This suggests that embeddings can not distinguish between domains in a mathematical sense.
6.2 Output Space: High-density “Pattern Collapse
The output space in mathematical reasoning exhibits a unique high-density characteristic, which we call “pattern collapse”. As shown in Figure 1, the distribution of output embeddings across various domains in the mathematical reasoning scenarios is significantly concentrated and indistinguishable.
Two reasons cause this phenomenon: (1) Expression level: Output space of mathematical reasoning is scalar [37], resulting in a compressed search space with a higher probability of overlap in two widely divergent questions. For example, “$1+3=$” and “$\int_{1}^{3}x\mathrm{d}x=$” are both “$4$”; (2) Token level: We usually categorize different mathematical reasoning domains from a mathematical sense. However, GLMs model real numbers or mathematical expressions not in a mathematical sense, but based on the average embedding of all discrete token embeddings after sequence tokenization. Through tokenization, two expressions that are very different in the mathematical sense (e.g., Two numbers that are far apart on the real number line) may share many tokens because regular mathematical expressions are only taken from 0-9 number tokens and a limited number of special symbols, such as decimal points, slashes, and root signs. Thus, the collapse occurs at the token level during the autoregressive prediction of each token.
Table 6: Statistics about output tokens on mathematical reasoning (seven different task types) and text generation (translation and summarization).
Task Type $N$ $N_{\mathcal{T}}$ $D$ $C$
Mathematical Reasoning
Arithmetic(easy) 16136 14 99.9% 0.04%
Arithmetic(hard) 5663 16 99.7% 0.05%
Algebra 5234 107 98.0% 0.33%
Geometry 2615 75 97.1% 0.23%
Cnt.&Prob. 2524 43 98.3% 0.13%
Num.Theory 2395 71 97.1% 0.22%
Precalculus 3388 84 97.5% 0.26%
Average 5422 58 98.9% 0.18%
Text Generation
Translation 2500 1065 57.4% 3.32%
5000 1832 63.3% 5.10%
10000 2980 70.2% 9.31%
Average 5833 1959 66.4% 6.12%
Summarization 2500 1265 49.4% 4.01%
5000 1970 60.6% 6.16%
10000 3192 68.0% 9.98%
Average 5833 2142 63.2% 6.69%
We conduct the following statistical analysis to demonstrate the universality of “pattern collapse” across various mathematical reasoning tasks: We categorize the mathematical tasks into various types across different domains and difficulties. In each task, we count the token number $N$ and the token type number $N_{\mathcal{T}}$ in the dataset, then compute the token duplication rate $D$ and the vocab coverage $C$ as follows: We use Llama-2 tokenizator [50], whose token type number in vocab $N_{\mathcal{V}}$ is 32000. The computation metric is:
$D=1-\frac{N_{\mathcal{T}}}{N},~{}~{}~{}C=\frac{N_{\mathcal{T}}}{N_{\mathcal{V}}}.$ (8)
We also test translation and summarization tasks by taking samples with the same token size as the mathematical reasoning dataset for a clear comparison. Table 6 presents the statistics data, we find that: (i) The average token duplication rate was 98.9% on all math tasks, and even a staggering 99.9% on some easy arithmetic tasks; In contrast, the token duplication rate on the text generation task is only about 60%, with about 2000 token types, and still increasing as the total number of tokens increases. These data and comparisons demonstrate that pattern collapse occurs on mathematical reasoning and not on text generation. (ii) Token repetition rate exceeded 97% on all seven math tasks of different difficulties and types. From these conclusions, we can demonstrate that the “pattern collapse” occurs on generally all types of mathematical reasoning tasks.
6.3 Can Chain-of-Thought Address “pattern collapse”?
A straightforward approach to addressing “pattern collapse” in the output space is to leverage chain-of-thought (CoT) techniques [54, 17, 62, 52, 61] to expand the output space size. Likewise, we adopt the solution steps associated with each sample in the MATH dataset as the output and employ SimCSE to derive embedding representations. The experimental setup aligns with Section 4.1, and the results are shown in Table 5(b). We note a similar phenomenon as in Table 5(a), i.e., the detection accuracy under different ID-OOD pairs varies greatly, and thus the detection randomness is more pronounced. This suggests that despite that CoT expands the output space size, the output answer is still essentially related to the difficulty and digit of mathematical reasoning, and the semantic embedding representation cannot reflect these features accurately.
7 Conclusion
Limitations
Our limitation mainly lies in the relatively small sizes of datasets used in our experiments. Due to the difficulty of collecting and labeling mathematical reasoning data, dataset sizes in this field are generally small, mostly in the hundreds or thousands, making it difficult to obtain millions of data for training and reasoning as in translation or summarization tasks. To address this, we adopt test sampling to reduce the randomness under small-scale testing and mitigate the data imbalance.
Ethics Statement
Acknowledgment
This work is supported by the Alibaba Innovative Research Program. Yiming and Rui are partially supported by the General Program of the National Natural Science Foundation of China (62176153), the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), and SMP-Zhipu.AI Large Model Cross-Disciplinary Fund, the Science and Technology Development Fund. Zhuosheng is partially supported by the National Natural Science Foundation of China (62406188). This work is also partially supported by the Science and Technology Development Fund, Macau SAR (Grant Nos. FDCT/060/2022/AFJ, FDCT/0070/2022/AMJ), and the Multi-year Research Grant from the University of Macau (Grant No. MYRG-GRG2023-00006-FST-UMDF).
Appendix
Appendix A Related Work
@@
Appendix B Motivation Visualization Details of Figure 1
B.1 Projection Setting
B.2 More Examples
Input Output
Domain 1: Algebra
Domain 2: Geometry
Domain 3: Number Theory
Domain 4: Precalculus
Simplify $\frac{1}{1-\tan^{2}x}+\frac{1}{1-\cot^{2}x}.$ 1
Input Output
Domain 1: Bible
Domain 2: News
Domain 3: TED
Domain 4: Health
Appendix C Theoretical Analysis of Section 2.1
@@ C.1 Problem Setup
For a given sample $s_{i}$, $[{\bm{y}_{i}}]_{l}\in\mathbb{R}^{d}(1\leq l\leq L)$ denote the hidden embedding of the $l$-th layer. We define the embedding coordinate in the latent space as $\{(l,[{\bm{y}_{i}}]_{l})\}_{l=1}^{L}$.
@@
We assume the ${\bm{F}}_{i}(x)=\left[F_{i1}(x),F_{i2}(x),...,F_{id}(x)\right]^{\top}:\mathbb{R}\rightarrow\mathbb{R}^{d}$ with $d$ independent components fits the $L$ coordinates, representing a continuous learning trajectory for $s_{i}$, so $[{\bm{y}_{i}}]_{l}={\bm{F}}_{i}(l)$. ${\bm{F}}_{i}(x)$ is taken from the functional space $\bm{X}$, so $\{{\bm{F}}_{i}(l)\}_{l=1}^{L}$ are all independent variables. We constrain each component function of this ${\bm{F}}_{i}(x)$ to satisfy the $m$-order $(m\geq 3)$ derivability property. Under this setting, the definition of dimension-independent trajectory volatility (Eq. 2) equates to
${\bm{V}}(s_{i})$ $=\left[V_{1}(s_{i}),V_{2}(s_{i}),...,V_{L}(s_{i})\right]^{\top}$ (9)
@@
We specify that $[{\bm{y}}_{i}]_{L}={\bm{F}}_{i}(L)\sim\mathcal{N}(\bm{c},\Sigma^{2})$, where
$\bm{c}=[c_{1},c_{2},...,c_{d}]^{\top},~{}~{}~{}\Sigma=\mathrm{diag}(\delta_{1},\delta_{2},\cdots,\delta_{d}).$ (10)
According to the pattern collapse property under different tasks, we can constraint the endpoint embedding ${\bm{F}}_{i}(L)$ in the output space:
  • For mathematical reasoning with pattern collapse, $\Sigma\rightarrow\bm{O}$, so we approximate that ${\bm{F}}_{i}(L)\equiv\bm{c}$;
  • For text generation without pattern collapse, $\Sigma\neq\bm{O}$, so ${\bm{F}}_{i}(L)\not\equiv\bm{c}$.
@@
Theorem C.1 (Main Theorem)
@@
$\mathbb{E}_{\{{\bm{F}}_{i}(l)\}_{l=1}^{L},~{}\{{\bm{F}}_{j}(l)\}_{l=1}^{L}~{}\sim~{}\mathcal{U}(\mathbb{R}^{d})}\left\{{\bm{V}}(s_{i})-{\bm{V}}(s_{j})\neq{\bm{0}}|{\bm{F}}_{i}(L),{\bm{F}}_{j}(L)\equiv\bm{c}\right\}$
@@ C.2 Prelinimary
Proposition C.1 (Lagrange Remainder Term)
@@
Lemma C.1 (Error Bound for the Midpoint Rule)
@@
Following Theorem C.1, we do the inequality scaling as below:
@@
Lemma C.2 (Differential-Integral Error Order Estimation)
@@
Following Lemma C.1, we do the inequality scaling for the first-order derivative summation:
@@
C.3 Main Theorem Proof
Finally, we formally prove our main theorem C.1 in this sub-section, thereby verifying the Hypotheses 1 in the main paper. Since the $d$ components of ${\bm{V}}(s_{i})-{\bm{V}}(s_{j})$ are independent of each other, we only consider the $d$-th dimensional component and the rest of the components $V_{d}(s_{i})-V_{d}(s_{j})$ can be proved in the same way.
Proof.  According to Lemma C.2, we can approximate $V_{d}(s_{i})-V_{d}(s_{j})$ as
@@ C.3 Main Theorem Proof
We assume that the first interval $(1,i_{1})$ and $(1,j_{1})$ are constant positive intervals (the same can be proven for constant negative intervals). In this setting, we can continue to simplify Eq.22:
@@
$\bm{c}=\left[(-1)^{i_{p}-1}\cdot F_{id}(L)-(-1)^{j_{q}-1}\cdot F_{jd}(L)\right],$ (26)
then we can rewrite Eq.25 to the matrix form:
$\int_{1}^{L}\left|\frac{\mathrm{d}F_{id}}{\mathrm{d}x}\right|~{}\mathrm{d}x-\int_{1}^{L}\left|\frac{\mathrm{d}F_{jd}}{\mathrm{d}x}\right|~{}\mathrm{d}x=\bm{A}\bm{x}+\bm{c},$ (27)
where $\bm{A}$ is the coefficient matrix, and $\bm{x}$ is the unknown variable vector:
$\bm{A}=[-1,1,\underbrace{2,-2,2,-2,...,(-1)^{p-1}\cdot 2}_{p~{}\text{dimensions}},\underbrace{-2,2,-2,2,...,(-1)^{q}\cdot 2}_{q~{}\text{dimensions}}]\in\mathbb{R}^{1\times(p+q+2)}$ (28)
$\bm{x}=[F_{id}(1),F_{jd}(1),F_{id}(i_{1}),...,F_{id}(i_{p}),F_{jd}(j_{1}),...,F_{jd}(j_{q})]^{\top}\in\mathbb{R}^{(p+q+2)\times 1}$
We let $i_{p}\equiv j_{q}(\mathrm{mod}~{}2)$, and return to the condition in the mathematical expectation on both sides of the inequality in the main theorem C.1 for the following categorical discussion:
  • If $F_{id}(L)=F_{jd}(L)=c_{d}$, then $\bm{c}=\bm{0}$, so $V_{d}(s_{i})-V_{d}(s_{j})=0\Longrightarrow\bm{Ax}=\bm{0}$;
  • If $F_{id}(L),F_{jd}(L)\sim\mathcal{N}(c_{d},\delta_{d}^{2})$, then $\bm{c}\not\equiv\bm{0}$, so $V_{d}(s_{i})-V_{d}(s_{j})\neq 0\Longrightarrow\bm{Ax}=-\bm{c}$.
We denote $N(A)$ as the solution space of $\{\bm{x}|\bm{Ax}=\bm{0}\}$ and $P(A)$ as the solution space of $\{\bm{x}|\bm{Ax}=-\bm{c}\}$. For $x\in P(A)$, $x=x_{p}$ or $x=x_{p}+x_{n}$ where $x_{p}$ is the special solution and $x_{n}$ is the solution in zero space, i.e., $x_{n}\in N(A)$. Since $\mathrm{rank}(A)=1<p+q+2$, the special solution $x_{p}$ exists, so the size of the solution space $P(A)$ is larger than $N(A)$. In this case, when the variable $\bm{x}$ is sampled from the real number space, its probability of being in $N(A)$ is smaller, namely its probability of not being in $N(A)$ is larger. This is exactly equivalent to the form of the proof of Main Theorem C.1, so the proof is complete. $\square$
C.4 Extended Conclusion